OpenStudy (anonymous):

Rusty's hair grows at the rate of 1/4 inch per month. How many months will it take Rusty's hair to grow 5/8 inch? Explain your answer using expressions and words.

OpenStudy (anonymous):

First get a common denominator

OpenStudy (anonymous):

1/4 = x/8?

OpenStudy (anonymous):

Hello? Trying to help here but I can't just give you answer. You need to participate.

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

Nope. We multiply by 2 on the bottom so we multiply by 2 on the top. Try again

OpenStudy (anonymous):

1/4 = 2/8 OK? Do you see how i did that?

OpenStudy (anonymous):

1/4 = 2/8 so what most likley what you did was you multiply to the top 2 so it can go exact with the bottom gettting the answer by divide 1/4 / 2/8 which is 1/2?

OpenStudy (anonymous):

NOOOOO: Ok so now it takes one month for his hair to grow 2/8 inches. We need to find how many months it takes to grow 5/8. 2/8 *2 = 4/8. So we need 1/8 more. It has been two months so It will take 2 and a 1/2 months to grow 5/8 inches (answer) got it?

OpenStudy (anonymous):

Medal please I will give you one

OpenStudy (anonymous):

btw to get a common denominator you just multiply the same for the the numerator and the denominator. Example: 3/9 = x/18 9*2 = 18 3*2 = 6 New fraction: 3/9 = 6/18

OpenStudy (anonymous):

Hope this helped. How old are you, you may need a tutor - this is pretty basic stuff. If you need any more help just tag me.

OpenStudy (anonymous):

well that makes it more easier with the common denominator example, Oh okay! I understand now Thank you :)

OpenStudy (anonymous):

but thanks @twistnflip